Formic acid, HCHO2, burns in oxygen to form carbon dioxide and water as follows: HCHO2 (aq) + O2 (g) → 2 CO2 (g) + 2 H2O (l). If a 3.15-g sample of formic acid was burned in 2.0 L of oxygen, what volume of carbon dioxide would be produced? (Assume the reaction occurs at standard temperature and pressure, STP.

1
Expert's answer
2022-04-16T04:09:27-0400

"n(HCHO_2)=\dfrac{Mass}{Molar\ mass}=\dfrac{3.15\ g}{46\ g/mol}=0.06848\ mol"


At STP, one mole of gas occupies 22.4 L of volume (molar volume).


"n(O_2)=\dfrac{Volume}{Molar\ volume}=\dfrac{2.0\ L}{22.4\ L/mol}=0.08929\ mol"


Balanced equation: 2HCHO2 (aq) + O2 (g)2CO2 (g) + 2H2O (l)


Formic acid is the limiting reactant.


"n(CO_2)=n(HCHO_2)=0.06848\ mol"


"V(CO_2)=Amount*Molar\ volume=0.06848\ mol\ *\ 22.4\ L/mol =1.534\ L"